OpenStudy (anonymous):

Another integration: int ((1/(1+sin z + cos z))) dz

sam (.sam.):

Hmm looks like we need weierstrass for this? If you let \[u=\tan(\frac{z}{2}) \\ \\ du=\frac{1}{2}\sec^2(\frac{z}{2})dz\] At the same time using weierstrass subs you get \[\Large \int\limits \frac{2}{ \left(\frac{2 u}{u^2+1}+\frac{1-u^2}{u^2+1}+1\right)\left(u^2+1\right)} \, du\] Simplify \[\int\limits \frac{1}{1+u} \, du\] Solve

OpenStudy (anonymous):

@.Sam. What is weierstrass

OpenStudy (anonymous):

thanks , never read that .